# A Symmetry-Preserving Reduced-Order Observer

### v1

@@ Abstract
A symmetry-preserving, reduced-order state observer is presented for the unmeasured part of a system’s state, where the nonlinear system dynamics exhibit symmetry under the action of a Lie group. The proposed observer takes advantage of this symmetry through the use of a moving frame that constructs invariant mappings of the measurements. Sufficient conditions for the observer to be asymptotically stable are developed by studying the stability of an invariant error system. As an illustrative example, the observer is applied to the problem of rigid-body velocity estimation, which demonstrate how exploiting the symmetry of the system can simplify the stabilization of the estimation error dynamics.
@@ I INTRODUCTION
Methods for designing state observers for nonlinear systems are limited and there are no general techniques that guarantee global convergence of the estimation error, as there are in the linear case [1, Ch. 15]. Provably effective state estimation strategies are inevitably limited to special classes of systems. Here, we leverage symmetries in a dynamical system’s structure to aid observer design and stability analysis. These so-called symmetry-preserving observers presented by Bonnabel et al. in [2] and [3] are adapted to systems whose dynamics are invariant under the action of a Lie group. The idea is to design an observer that is also invariant, i.e., for which the observer dynamics also preserve this Lie group symmetry. This approach allows the observer’s convergence properties to be analyzed more easily because of simplifications afforded by the system symmetry.
@@
Existing approaches to symmetry-preserving observers only consider the full-order case, however, in which the entire state of the system is estimated ([2, 3, 4]). In many scenarios, part of the system’s state may be known with negligible error or may be obtained as the output of an observer whose design is independent of the rest of the system’s state. For example, attitude observers for aircraft or spacecraft often do not rely on the rigid body’s translational dynamics ([5, 6]). Another example is the problem of wind estimation from aircraft motion ([7, 8, 9]), where the main goal is to obtain estimates of wind and air-relative velocity – not to re-estimate the aircraft’s position, attitude, and angular velocity. More generally, the problem of disturbance estimation falls into this category where the internal state of the system is known but the disturbance is not. In these scenarios, reduced-order observers, in the sense of Karagiannis et al. in [10] and [11], are of particular interest where only the unmeasured part of the system’s state is estimated. The aim of observer design is to render a particular set, characterized by zero state estimation error, positively invariant and globally asymptotically attractive. Here, we focus on reduced-order observers that are also symmetry-preserving.
@@
The remainder of this paper is organized as follows. Section II introduces the preliminary concepts that will be used in the development of a reduced-order, symmetry-preserving pre-observer in Section III. Next, sufficient conditions for the pre-observer to be an asymptotically stable observer are presented in Section IV. Finally, the main results are applied to the example of rigid-body velocity estimation in Section V followed by concluding remarks in Section VI.
@@ II-A2 The Moving Frame [2, 13, 14]
A moving frame is a mapping $\gamma:\mathcal{X}\to G$ that has the following equivariance property (illustated in Fig. 1):
@@
It may be associated with a coordinate cross-section $\mathcal{K}$ that transversely intersects $g$-orbits on $\mathcal{X}$. Informally, for an $r$-dimensional Lie group $G$ acting freely on the $n$-dimensional manifold $\mathcal{X}$, let $\varphi_{g}^{\mathrm{inv}}$ be the part of $\varphi_{g}$ that maps to an $r$-dimensional submanifold of $\mathcal{X}$ such that it is invertible with respect to $g$ in a neighborhood of the identity element $e\in G$. Then, one can select a constant ${k}$ in the image of $\varphi_{g}^{\mathrm{inv}}$ that defines the unique point at which the $g$-orbit of a generic point ${x}$ intersects the $(n-r)$-dimensional cross-section $\mathcal{K}$. In other words, it is obtained by solving the normalization equation
@@ II-A3 Invariant Dynamics [2, 12]
where ${x}(t)\in\mathcal{X}$ (a differentiable manifold), ${u}(t)\in\mathcal{U}$ (a set), and ${f}(\cdot,{u}):\mathcal{X}\to\mathsf{T}_{x}\mathcal{X}$ (a vector field on $\mathcal{X}$ for each ${u}\in\mathcal{U}$). Here, $\mathsf{T}_{x}\mathcal{X}$ denotes the tangent space to $\mathcal{X}$ at ${x}$. Let $\{({\varphi}_{g}({x}),{\psi}_{g}({u}))\}_{g\in G}$ be a transformation group on $\mathcal{X}\times\mathcal{U}$. The mapping ${\varphi}_{g}:\mathcal{X}\to\mathcal{X}$ induces the tangent map $\mathrm{T}{\varphi}_{g}({x}):\mathsf{T}_{x}\mathcal{X}\to\mathsf{T}_{{\varphi}_{g}({x})}\mathcal{X}$ at ${x}$. Note if $\mathcal{X}=\mathbb{R}^{n}$, then $\mathrm{T}{\varphi}_{g}({x})$ is simply the Jacobian matrix, $\partial{\varphi}_{g}({x})/\partial{x}$. The system (2) is called $G$-invariant if
${f}({\varphi}_{g}({x}),{\psi}_{g}({u}))=\mathrm{T}{\varphi}_{g}({x})\cdot{f}({x},{u})$
Here, $\cdot$” denotes the application of the tangent map to a tangent vector. It follows that the tangent map of ${\varphi}_{g\ast h}({x})={\varphi}_{g}({\varphi}_{h}({x}))$ satisfies
$\mathrm{T}{\varphi}_{g\ast h}({x})=\mathrm{T}{\varphi}_{g}({\varphi}_{h}({x}))\cdot\mathrm{T}{\varphi}_{h}({x})$
A function $I:\mathcal{X}\times\mathcal{U}\to\mathbb{R}$ is called an invariant if $I({\varphi}_{g}({x}),{\psi}_{g}({u}))=I(x,u)$ for all $g\in G$. Suppose $G$ acts freely on $\mathcal{X}$. Then, there locally exist $n-r$ functionally independent invariants $(I_{1}({x}),\dots,I_{n-r}({x}))$.
@@ II-B Immersion and Invariance Observers [10, 11]
Consider a dynamical system whose state is described by a measured part, ${y}\in\mathcal{Y}\subseteq\mathbb{R}^{p}$, and an unmeasured part, ${x}\in\mathcal{X}\subseteq\mathbb{R}^{n}$, with dynamics
@@
defined by mappings ${\theta}$ and ${\phi}$ that are left invertible with respect to their first argument such that $\mathcal{Z}$ is positively invariant and (globally) asymptotically attractive. The estimate of ${x}$ is then given by
$\hat{{x}}={\phi}^{(\mathrm{L},\cdot)}({\theta}({z},{y}),{y})$
where ${\phi}^{(\mathrm{L},\cdot)}$ denotes the functional left inverse of ${\phi}$ with respect to its first argument (i.e., ${\phi}^{(\mathrm{L},\cdot)}({\phi}({x},{y}),{y})={x}$).
@@ III INVARIANT PRE-OBSERVER
Consider a system whose state is given by a measured part, ${y}\in\mathcal{Y}$, and an unmeasured part, ${x}\in\mathcal{X}\subseteq\mathbb{R}^{n}$. Here, $\mathcal{Y}$ is a $p$-dimensional differentiable manifold and $\mathcal{X}$ is an open subset of $\mathbb{R}^{n}$ that contains the origin. The dynamics of this system are given by
@@
where ${u}\in\mathcal{U}$ is the known “input” to the system. It is not necessarily just composed of control inputs, but rather is a known signal on which a particular Lie group acts. Here, the dynamics of the measured part of the state, ${y}$, may be expressed intrinsically, that is, without specifying a local coordinate chart.
We consider systems of the form (4) that are invariant under actions of some Lie group, $G$.
@@
$\mathrm{T}{\varphi}_{g}({x})\cdot{f}({x},{y},{u})$ $={f}({\varphi}_{g}({x}),{\varrho}_{g}({y}),{\psi}_{g}({u}))$
$\mathrm{T}{\varrho}_{g}({y})\cdot{h}({x},{y},{u})$ $={h}({\varphi}_{g}({x}),{\varrho}_{g}({y}),{\psi}_{g}({u}))$
@@
We can now describe what it means for a reduced-order pre-observer in the sense of [10, 11] to be symmetry-preserving under the transformation group considered in Assumption 1. Briefly, a pre-observer is an observer for which there is not (yet) any claim about error convergence. We postulate the form of an observer for the unmeasured part of the state, ${x}$, that preserves invariance of the state estimate dynamics. Inspired by [2], consider the following definition.
@@
$\dot{\hat{{x}}}={\alpha}(\hat{{x}}-{\beta}({y}),{y},{u})+\mathrm{T}{\beta}({y})\cdot{h}({x},{y},{u})$ (7)
@@
This prescription of the zero-error manifold $\mathcal{Z}$ is not quite as general as the case described in [11]. We instead consider the condition that defines $\mathcal{Z}$ to be linear in ${z}$ and ${x}$ (and for ${z}$ to be the same dimension as ${x}$). In other words, we choose ${\theta}({z},{y})={z}$ and ${\phi}({x},{y})={x}-{\beta}({y})$ in Eq. (3). This simplification reveals an intuitive choice for ${\beta}$ in the following lemma based on the underlying geometry.
@@
$\mathrm{T}{\beta}({\varrho}_{g}({y}))\cdot\mathrm{T}{\varrho}_{g}({y})=\mathrm{T}{\varphi}_{g}({\beta}({y}))\cdot\mathrm{T}{\beta}({y})$ (11)
@@
${\alpha}({z},{y},{u})={f}({z}+{\beta}({y}),{y},{u})-\mathrm{T}{\beta}({y})\cdot{h}({z}+{\beta}({y}),{y},{u})$ (12)
Then, the dynamical system
$\dot{{z}}={\alpha}({z},{y},{u})$ (13)
with output
$\hat{{x}}={z}+{\beta}({y})$ (14)
is a $G$-invariant, reduced-order pre-observer.
@@
${F}(\hat{{x}},{x},{y},{u})={\alpha}(\hat{{x}}-{\beta}({y}),{y},{u})+\mathrm{T}{\beta}({y})\cdot{h}({x},{y},{u})$
@@
$\mathrm{T}{\varphi}_{g}(\hat{{x}})\cdot{F}(\hat{{x}},{x},{y},{u})=\mathrm{T}{\varphi}_{g}(\hat{{x}})\cdot{f}(\hat{{x}},{y},{u})\\ -\mathrm{T}{\varphi}_{g}(\hat{{x}})\cdot\mathrm{T}{\beta}({y})\cdot\left({h}(\hat{{x}},{y},{u})-{h}({x},{y},{u})\right)$
The assumed linearity of ${\varphi}_{g}$ implies $\mathrm{T}{\varphi}_{g}({x})$ does not depend on the choice of base point ${x}$. Therefore, $\mathrm{T}{\varphi}_{g}(\hat{{x}})=\mathrm{T}{\varphi}_{g}({\beta}({y}))$, and Lemma 1 can be used along with the invariance of ${f}$ to obtain
$\mathrm{T}{\varphi}_{g}(\hat{{x}})\cdot{F}(\hat{{x}},{x},{y},{u})={f}({\varphi}_{g}(\hat{{x}}),{\varrho}_{g}({y}),{\psi}_{g}({u}))\\ -\mathrm{T}{\beta}\left({\varrho}_{g}({y})\right))\cdot\mathrm{T}{\varrho}_{g}({y})\cdot\left({h}(\hat{{x}},{y},{u})-{h}({x},{y},{u})\right)$
@@
$\mathrm{T}{\varphi}_{g}(\hat{{x}})\cdot{F}(\hat{{x}},{x},{y},{u})={f}({\varphi}_{g}(\hat{{x}}),{\varrho}_{g}({y}),{\psi}_{g}({u}))\\ -\mathrm{T}{\beta}\left({\varrho}_{g}({y})\right)\cdot\Big{(}{h}({\varphi}_{g}(\hat{{x}}),{\varrho}_{g}({y}),{\psi}_{g}({u}))\\ -{h}({\varphi}_{g}({x}),{\varrho}_{g}({y}),{\psi}_{g}({u}))\Big{)}$
@@
$\mathrm{T}{\varphi}_{g}(\hat{{x}})\cdot{F}(\hat{{x}},{x},{y},{u})={\alpha}({\varphi}_{g}(\hat{{x}})-{\beta}({\varrho}_{g}({y})),{\varrho}_{g}({y}),{\psi}_{g}({u}))\\ +\mathrm{T}{\beta}\left({\varrho}_{g}({y})\right)\cdot{h}({\varphi}_{g}({x}),{\varrho}_{g}({y}),{\psi}_{g}({u}))$
@@
$\mathrm{T}{\varphi}_{g}(\hat{{x}})\cdot{F}(\hat{{x}},{x},{y},{u})={F}({\varphi}_{g}(\hat{{x}}),{\varphi}_{g}({x}),{\varrho}_{g}({y}),{\psi}_{g}({u}))$
That is, the system (12)-(14) is $G$-invariant. Next, we show the zero error manifold $\mathcal{Z}$ given in (8) is positively invariant. Since ${z}-{x}+{\beta}({y})={0}$ on $\mathcal{Z}$, we verify that
${\alpha}({x}-{\beta}({y}),{y},{u})-{f}({x},{y},{u})+\mathrm{T}{\beta}({y})\cdot{h}({x},{y},{u})$
$\quad={f}({x},{y},{u})-\mathrm{T}{\beta}({y})\cdot{h}({x},{y},{u})$
$\qquad-{f}({x},{y},{u})+\mathrm{T}{\beta}({y})\cdot{h}({x},{y},{u})$
$\quad={0}$
Thus, referring to (13), trajectories originating in $\mathcal{Z}$ remain in $\mathcal{Z}$. It follows that (12)-(14) is a $G$-invariant reduced-order pre-observer. ∎

### v2

@@ Abstract
A symmetry-preserving, reduced-order state observer is presented for the unmeasured part of a system’s state, where the nonlinear system dynamics exhibit symmetry under the action of a Lie group. Leveraging this symmetry with a moving frame, the observer dynamics are constructed such that they are invariant under the Lie group’s action. Sufficient conditions for the observer to be asymptotically stable are developed by studying the stability of an invariant error system. As an illustrative example, the observer is applied to the problem of rigid-body velocity estimation, which demonstrates how exploiting the symmetry of the system can simplify the stabilization of the estimation error dynamics.
@@ I INTRODUCTION
Methods for designing state observers for nonlinear systems are limited, and there are no general techniques that guarantee global convergence of the estimation error as there are in the linear case [1, Ch. 15]. Provably effective state estimation strategies are inevitably limited to special classes of systems, motivating considerable attention in the past several decades to nonlinear observers. Early results generally rely either on Lipschitz conditions (e.g., [2, 3, 4]) or finding a transformation of the system to some canonical form (e.g., [5, 6, 7]). For some systems, however, these methods are overly-conservative or difficult to implement. Passivity-based observers ([8, 9]) constitute an approach that aims to overcome these deficiencies by viewing observer design as output feedback control of a state estimation error system — a perspective inherently shared among nonlinear observer techniques as a result of the duality between nonlinear controllability and observability [10, Ch. 3].
Approaches to observer design that leverage the structure of the system are of particular interest. Specifically, the role of differential geometry in observer design has been explored, in which symmetries of a nonlinear dynamical system are preserved in the state observer and its state estimation error dynamics, aiding in design and stability analysis. From a Lagrangian perspective, Aghannan and Rouchon [11] leveraged the symmetry inherent in the coordinate-free Euler-Lagrange equations to design a nonlinear observer given measurements of the system’s configuration. This idea of leveraging symmetries in the dynamics was generalized by Bonnabel et al. to include nonlinear systems under a Lie group’s action [12] as well as systems defined on Lie groups [13]. The idea of these so-called symmetry-preserving observers is to design an observer that is also invariant, i.e., for which the observer dynamics also preserve this Lie group symmetry. This approach allows the observer’s convergence properties to be analyzed more easily because of simplifications afforded by the system symmetry. The invariant extended Kalman filter (EKF) developed in Bonnabel et al. [14] is a constructive, local symmetry-preserving observer that is also applicable to systems with measurement and process noise. The invariant EKF was shown by Barrau and Bonnabel in [15] to be a stable observer, admitting an exact linearization for group affine systems. From a preservative of equivariance, Mahony et al. [16] developed a nonlinear observer for kinematic systems with complete symmetry — that is, systems defined on homogeneous spaces. This work was generalized with the equivariant Kalman filter, which applies to general equivariant systems (see van Goor et al. [17] and references therein).
@@
Existing approaches to symmetry-preserving observers only consider the full-order case, however, in which the entire state of the system is estimated. In many scenarios, part of the system’s state may be known with negligible error or may be obtained as the output of an observer whose design is independent of the rest of the system’s state. For example, attitude observers for aircraft or spacecraft often do not rely on the rigid body’s translational dynamics (e.g., [18] and [19]). Another example is the problem of wind estimation from aircraft motion (e.g., [20, 21]), where the main goal is to obtain estimates of wind and air-relative velocity — not to re-estimate the aircraft’s position, attitude, and angular velocity. Estimating wind can be framed as a disturbance estimation problem, falling into the general category where the internal state of the system is known but the disturbance is not. Disturbance observers generally assume the dynamics of the disturbance are unknown (e.g., [22, 23]) and can also be applied to systems defined on Lie groups (e.g., [24, 25]). However, when the disturbance dynamics are known, they may be leveraged to obtain improved performance and/or stronger guarantees.
In scenarios where full-order observers are either unnecessary or impractical, reduced-order observers, in the sense of Karagiannis et al. in [26] and [27], are of particular interest in which only the unmeasured part of the system’s state is estimated. The aim of reduced-order observer design is to render a particular set, characterized by zero state estimation error, positively invariant and globally asymptotically attractive. However, the definitions of this set and the observer dynamics are non-constructive (although constructive approaches are available for certain classes of systems [9]). In this paper, we develop a reduced-order observer that is also symmetry-preserving. That is, the zero-error set and the observer dynamics are constructed in a way such that they are invariant under the Lie group’s action. Leveraging this symmetry leads to an invariant state estimation error system, simplifying observer parameter selection and stability analysis.
@@
The remainder of this paper is organized as follows. Section II introduces the preliminary concepts that will be used in the development of the pre-observer in Section III. Next, sufficient conditions for the pre-observer to be an asymptotically stable observer are presented in Section IV. Finally, the main results are applied to the example of rigid-body velocity estimation in Section V, followed by concluding remarks in Section VI.
@@ II-A2 The Moving Frame [12, 29, 30]
A moving frame is a mapping $\gamma:\mathcal{X}\to G$ that has the following equivariance property (illustrated in Fig. 1):
@@
It may be associated with a coordinate cross-section $\mathcal{K}$ that transversely intersects $G$-orbits on $\mathcal{X}$. Informally, for an $r$-dimensional Lie group $G$ acting freely on the $n$-dimensional manifold $\mathcal{X}$, let $\varphi_{g}^{\mathrm{inv}}$ be the part of $\varphi_{g}$ that maps to an $r$-dimensional submanifold of $\mathcal{X}$ such that it is invertible with respect to $g$ in a neighborhood of the identity element $e\in G$. Then, one can select a constant k{k} in the image of $\varphi_{g}^{\mathrm{inv}}$ that defines the unique point at which the $G$-orbit of a generic point $x$ intersects the (nr)(n-r)-dimensional cross-section $\mathcal{K}$. In other words, the moving frame can be obtained by solving the normalization equation
@@ II-A3 Invariant Dynamics [12, 28]
where $x(t)\in\mathcal{X}$ (a differentiable manifold), $u(t)\in\mathcal{U}$ (a set), and $f(\cdot,u):\mathcal{X}\to\mathsf{T}_{x}\mathcal{X}$ (a vector field on $\mathcal{X}$ for each $u\in\mathcal{U}$). Here, $\mathsf{T}_{x}\mathcal{X}$ denotes the tangent space to $\mathcal{X}$ at $x$. Let $\{(\varphi_{g}(x),\psi_{g}(u))\}_{g\in G}$ be a transformation group on $\mathcal{X}\times\mathcal{U}$, where $G$ is an $r$-dimensional Lie group. The mapping $\varphi_{g}:\mathcal{X}\to\mathcal{X}$ induces the tangent map $\mathrm{T}_{x}\varphi_{g}:\mathsf{T}_{x}\mathcal{X}\to\mathsf{T}_{\varphi_{g}(x)}\mathcal{X}$ at $x$. Note that if $\mathcal{X}=\mathbb{R}^{n}$, then $\mathrm{T}_{x}\varphi_{g}$ is simply the Jacobian matrix, $\partial\varphi_{g}(x)/\partial x$. The system (2) is called $G$-invariant if
$f(\varphi_{g}(x),\psi_{g}(u))=\mathrm{T}_{x}\varphi_{g}\big{(}f(x,u)\big{)}$
It follows that the tangent map of $\varphi_{g\ast h}(x)=\varphi_{g}(\varphi_{h}(x))$ satisfies
$\mathrm{T}_{x}\varphi_{g\ast h}=\mathrm{T}_{\varphi_{h}(x)}\varphi_{g}\circ\mathrm{T}_{x}\varphi_{h}$
A function $I:\mathcal{X}\times\mathcal{U}\to\mathbb{R}$ is called an invariant if $I(\varphi_{g}(x),\psi_{g}(u))=I(x,u)$ for all $g\in G$. Suppose $G$ acts freely on $\mathcal{X}$. Then, there locally exist nrn-r functionally independent invariants $(I_{1}(x),\dots,I_{n-r}(x))$.
@@ II-B Immersion and Invariance Observers [26, 27]
Consider a dynamical system whose state is described by an unmeasured part, $x\in\mathcal{X}\subset\mathbb{R}^{n}$, and a measured part, $y\in\mathcal{Y}\subset\mathbb{R}^{p}$, with dynamics
@@
defined by smooth mappings $\theta$ and $\phi$ that are left invertible with respect to their first argument such that $\mathcal{Z}$ is positively invariant and (globally) asymptotically attractive. The estimate of $x$ is then given by
$\hat{x}=\phi^{(\mathrm{L},\cdot)}(\theta(z,y),y)$ (5)
where $\phi^{(\mathrm{L},\cdot)}$ denotes the functional left inverse of $\phi$} with respect to its first argument; that is, $\phi^{(\mathrm{L},\cdot)}(\phi(x,y),y)=x$.
@@ III INVARIANT PRE-OBSERVER
Consider a system whose state is given by an unmeasured part $x\in\mathcal{X}\subseteq\mathbb{R}^{n}$ and a measured part $y\in\mathcal{Y}$. Here, $\mathcal{Y}$ is a $p$-dimensional differentiable manifold and $\mathcal{X}$ is an open subset of $\mathbb{R}^{n}$ that contains the origin. The dynamics of this system are given by
@@
where $u\in\mathcal{U}$ is the known “input” to the system. It is not necessarily just composed of control inputs; rather it is a known signal on which a particular Lie group acts. Here, the dynamics of the measured part of the state, $y$, may be expressed intrinsically, that is, without specifying a local coordinate chart.
We consider systems of the form (6) that are invariant under the action of some Lie group $G$.
@@
$\mathrm{T}_{x}\varphi_{g}\big{(}f(x,y,u)\big{)}$ $=f(\varphi_{g}(x),\varrho_{g}(y),\psi_{g}(u))$
$\mathrm{T}_{y}\varrho_{g}\big{(}h(x,y,u)\big{)}$ $=h(\varphi_{g}(x),\varrho_{g}(y),\psi_{g}(u))$
@@
We can now describe what it means for a reduced-order observer to be symmetry-preserving under the transformation group considered in Assumption 1. Briefly, a pre-observer is an observer for which there is not (yet) any claim about error convergence. We postulate a form for the observer of the unmeasured part of the state $x$ that preserves invariance of the state estimate dynamics. Inspired by [12], we propose the following definition.
@@
$\dot{\hat{x}}=\alpha(\hat{x}-\beta(y),y,u)+\mathrm{T}_{y}\beta\big{(}h(x,y,u)\big{)}$ (9)
@@
The prescription of the zero-error manifold $\mathcal{Z}$ here is not quite as general as the case described in [27]. We instead consider the condition that defines $\mathcal{Z}$ to be linear in $z$ and $x$ (and for $z$ to be the same dimension as $x$). In other words, we choose $\theta(z,y)=z$ and $\phi(x,y)=x-\beta(y)$ in (4). This simplification reveals an intuitive choice for the observer map $\beta$ in the following lemma based on the underlying geometry.
@@
$\mathrm{T}_{\varrho_{g}(y)}\beta\circ\mathrm{T}_{y}\varrho_{g}=\mathrm{T}_{\beta(y)}\varphi_{g}\circ\mathrm{T}_{y}\beta$ (13)
@@
$\alpha(z,y,u)=f(z+\beta(y),y,u)-\mathrm{T}_{y}\beta\big{(}h(z+\beta(y),y,u)\big{)}$ (14)
and let the observer map $\beta$ be given by (11). Then, the dynamical system (7) with output (8) is a $G$-invariant, reduced-order pre-observer.
@@
$F(\hat{x},x,y,u)=\alpha(\hat{x}-\beta(y),y,u)+\mathrm{T}_{y}\beta\big{(}h(x,y,u)\big{)}$
@@
$\mathrm{T}_{\hat{x}}\varphi_{g}\big{(}F(\hat{x},x,y,u)\big{)}=\mathrm{T}_{\hat{x}}\varphi_{g}\big{(}f(\hat{x},y,u)\big{)}\\ -\big{(}\mathrm{T}_{\hat{x}}\varphi_{g}\circ\mathrm{T}_{y}\beta\big{)}\big{(}h(\hat{x},y,u)-h(x,y,u)\big{)}$
The assumed linearity of $\varphi_{g}$ implies $\mathrm{T}_{x}\varphi_{g}$ does not depend on the choice of base point $x$. Therefore, $\mathrm{T}_{\hat{x}}\varphi_{g}=\mathrm{T}_{\beta(y)}\varphi_{g}$, and Lemma 1 can be used along with the invariance of $f$ to obtain
$\mathrm{T}_{\hat{x}}\varphi_{g}\big{(}F(\hat{x},x,y,u)\big{)}=f(\varphi_{g}(\hat{x}),\varrho_{g}(y),\psi_{g}(u))\\ -\big{(}\mathrm{T}_{\varrho_{g}(y)}\beta\circ\mathrm{T}_{y}\varrho_{g}\big{)}\big{(}h(\hat{x},y,u)-h(x,y,u)\big{)}$
@@
$\mathrm{T}_{\hat{x}}\varphi_{g}\big{(}F(\hat{x},x,y,u)\big{)}=f(\varphi_{g}(\hat{x}),\varrho_{g}(y),\psi_{g}(u))\\ -\mathrm{T}_{\varrho_{g}(y)}\beta\big{(}h(\varphi_{g}(\hat{x}),\varrho_{g}(y),\psi_{g}(u))\\ -h(\varphi_{g}(x),\varrho_{g}(y),\psi_{g}(u))\big{)}$
@@
$\mathrm{T}_{\hat{x}}\varphi_{g}\big{(}F(\hat{x},x,y,u)\big{)}=\alpha(\varphi_{g}(\hat{x})-\beta(\varrho_{g}(y)),\varrho_{g}(y),\psi_{g}(u))\\ +\mathrm{T}_{\varrho_{g}(y)}\beta\big{(}h(\varphi_{g}(x),\varrho_{g}(y),\psi_{g}(u))\big{)}$
@@
$\mathrm{T}_{\hat{x}}\varphi_{g}\big{(}F(\hat{x},x,y,u)\big{)}=F(\varphi_{g}(\hat{x}),\varphi_{g}(x),\varrho_{g}(y),\psi_{g}(u))$
That is, the system (9) is $G$-invariant. Next, we show the zero error manifold $\mathcal{Z}$ given in (10) is positively invariant. Since $z-x+\beta(y)=0$ on $\mathcal{Z}$, it is sufficient to verify that
$\alpha(x-\beta(y),y,u)-f(x,y,u)+\mathrm{T}_{y}\beta\big{(}h(x,y,u)\big{)}$
$\quad=f(x,y,u)-\mathrm{T}_{y}\beta\big{(}h(x,y,u)\big{)}$
$\qquad-f(x,y,u)+\mathrm{T}_{y}\beta\big{(}h(x,y,u)\big{)}=0$
Thus, referring to (7), trajectories originating in $\mathcal{Z}$ remain in $\mathcal{Z}$. It follows that (14)–(8) is a $G$-invariant reduced-order pre-observer. ∎
As an improvement over the general reduced-order observer described in §II-B, Theorem 1 leverages symmetry to construct a reduced-order pre-observer. The functions $\theta$, $\phi$, and $\alpha$ in (3)–(5) are formulated using the moving frame $\gamma$. The equivariance of the moving frame is the key property that makes this pre-observer $G$-invariant.